using zero reference potential at infinity, determine the amount by which point a charge of 8x10-8 c alter the electric potential at point 1.5m away where the charge is positive


1
Expert's answer
2022-03-13T18:46:13-0400

The electric potential

"V=k\frac{q}{r}=9*10^9*\frac{8*10^{-8}}{1.5}=480\:\rm V"